OpenStudy (anonymous):

help guys. Evaluate the composite function f(g(x)) for x = 31. f(x) = 9x + 1 g(x) =√x+5

OpenStudy (danjs):

is g(x) \[\sqrt{x+5}\]

OpenStudy (anonymous):

yes

OpenStudy (dan815):

put it into f(x)

OpenStudy (anonymous):

ok

OpenStudy (dan815):

f(g(x) = 9(g(x))+1 \[ =9\sqrt{x+5} +1\] x=31 \[ =9\sqrt{31+5} +1\] simplify.

OpenStudy (anonymous):

56.1?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

9*sqrt36+1 =9*6+1=55